# Suitability of KANs for Computer Vision: A preliminary investigation

### v1

@@ Abstract
Kolmogorov-Arnold Networks (KANs) introduce a paradigm of neural modeling that implements learnable functions on the edges of the networks, diverging from the traditional node-centric activations in neural networks. This work assesses the applicability and efficacy of KANs in visual modeling, focusing on the image recognition task. We mainly analyze the performance and efficiency of different network architectures built using KAN concepts along with conventional building blocks of convolutional and linear layers, enabling a comparative analysis with the conventional models. Our findings are aimed at contributing to understanding the potential of KANs in computer vision, highlighting both their strengths and areas for further research. Our evaluation [^1] This manuscript is a work in progress, which includes results from an ongoing exploration in this direction. The results and discussion are expected to evolve over time. This independent effort is intended to contributes to our understanding of KANs for visual modeling, and not to provide a yes/no answer to the question ‘if KANs should replace conventional neural modeling techniques for computer vision’. shows that whereas KAN-based architectures perform in-line with the original claims of [1] for performance and model-complexity in the case of simpler vision datasets like MNIST, the advantages seem to diminish even for slightly more complex datasets like CIFAR-10.
@@ I Introduction
Several implementations and research efforts are currently being made to explore the potential of KANs in various domains. The Convolutional-KANs project [24] extends the idea of KANs to convolutional layers, replacing the classic linear transformation of the convolution with learnable non-linear activations in each pixel. The Vision-KAN repository [25] explores the possibility of MLPs with KANs in Vision Transformers. The KAN-GPT project [26] implements Generative Pre-trained Transformers (GPTs) using KANs for language modeling tasks, demonstrating the potential of KANs in natural language processing applications. The KAN-GPT-2 repository [27] trains small GPT-2 style models using KANs instead of MLPs. The KANeRF project [28] integrates KANs into Neural Radiance Fields (NeRF) for view synthesis tasks. The KAN-UNet repository [29] implements a U-Net architecture with Kolmogorov-Arnold Convolutions (KA convolutions) for image segmentation tasks. Similarly, the kansformers project [30] explores the use of KANs in Transformer architectures, replacing the traditional linear layers with KAN layers.
Although there is currently a large interest in developing implementation repositories providing various KAN variants pertinent to visual modeling, a formal scientific effort focusing on the suitability of KANs in general and the emerging repositories in particular for visual modeling tasks is still missing. This work is aimed at partially addressing this gap as a preliminary effort focusing on providing independent empirical evidence of KANs’ suitability for vision in the context of image recognition task. We replicate ideas from Convolutional-KANs [24] and torch-conv-kan [31] implementations to evaluate their performance in terms of accuracy, training efficiency, and model parameters while experimenting with datasets such as MNIST and CIFAR. Our experiments are expected to provide insights into the strengths and limitations of KANs in handling complex visual data. The study aims at helping understand answers to the following questions:
@@ II-E Wavelet KANs (WavKANs)
WavKANConvs provide a powerful extension to KANs, incorporating wavelet-based frequency analysis into the adaptive filtering process of convolutional networks. This combination enables the extraction of complex patterns with enhanced localization in both time and frequency domains, potentially leading to superior performance in various visual tasks. The deeper and complex architectures are presented based on these adaptations, the detailed workings and configuration of KConvKANs is described in the next section.
@@ III Experiments
This section details the experimental setup and procedures used to evaluate the performance of networks on MNIST and CIFAR-10 datasets (in this version). The goal of these experiments is to assess the suitability of KANs for the fundamental computer vision task of recognition, focusing on aspects such as accuracy, model size, training time, and parameter efficiency. The architectural overview and the internal visualizations of a KConvKAN used in our experiments are illustrated in Fig. 3 for a clear understanding. The shown KConvKAN in Fig. 3(a) depicts the flow of data through the network that consists of two KAN Convolutional (KANConv) layers, each followed by a MaxPool2D layer, then a Flatten layer, and finally a KAN Linear layer that maps the flattened output of the convolutional layers to the final classification output via a log-softmax function. Fig.  3(b) provides insights into the feature maps and learned spline weights at different stages of the KConvKAN when processing a sample image. The feature map progression is shown after each KANConv and MaxPool2D layer. The 3D spline weights and heat map representation of spline weights in the KAN Linear layer are also visualized. The visualization is meant to demonstrate the transformation of the input through the network layers for understanding.
III-A Experimental Setup
@@ III-D Results
III-E Claims Validation
@@
III-E1 Claim-1: Accuracy
The original work on KANs [1] identifies that smaller KAN models can match or exceed the accuracy of larger MLPs in function fitting tasks. We can analyze this claim through our experiments on MNIST and CIFAR10 datasets. Our preliminary results show that KANs show competitive results with ConvNets of similar size. However, their advantage seems to diminish as task complexity increases.
III-E2 Claim-2: Computational Efficiency
@@
III-E3 Claim-3: Scalability
According to the KAN paper, these networks exhibit faster neural scaling laws than MLPs. However, when scaling up the networks to accommodate the high-dimensional data from CIFAR10, KANs required significant tuning of spline parameters to maintain performance, which adds to the training complexity. This observation leads to a nuanced view of the scalability claim, suggesting that practical scalability is contingent on the ability to manage increasing model complexity efficiently. The scalability of KANs to larger datasets, e.g., ImageNet, will be considered in the later versions.
@@ IV Conclusion
This work critically evaluates the efficacy of Kolmogorov-Arnold networks in the basic computer vision task of image recognition, contrasting their performance with conventional networks across MNIST and CIFAR-10 datasets. This work is a continuous effort that will be updated as more data and results are available. Our current findings reveal that while KANs do offer variation in the traditional neural network architectures, their superior accuracy and scalability are yet to be clearly evident for even slightly more complex vision datasets, e.g., CIFAR-10. The distinctive architecture of KANs, which integrates learnable spline functions directly onto the edges, provides an innovative approach to neural network design and opens new avenues for enhancing model interpretability and efficiency. The practical application of KANs in complex visual tasks require further optimization and exploration to fully harness their theoretical advantages. The research community is already engaged in leveraging KANs for vision or related tasks. Future work should focus on improving KAN-based architecture for scalability and generalization in diverse application scenarios within computer vision. Overall, whereas our investigation is yet to establish a clear performance advantage of KANs in vision domain, it does support the argument that investing in KAN-based vision models is worthwhile. With the large interest of vision and non-vision community in KANs, we can expect to see interesting developments in KANs for computer vision.

### v2

@@ Abstract
Kolmogorov-Arnold Networks (KANs) introduce a paradigm of neural modeling that implements learnable functions on the edges of the networks, diverging from the traditional node-centric activations in neural networks. This work assesses the applicability and efficacy of KANs in visual modeling, focusing on fundamental recognition and segmentation tasks. We mainly analyze the performance and efficiency of different network architectures built using KAN concepts along with conventional building blocks of convolutional and linear layers, enabling a comparative analysis with the conventional models. Our findings are aimed at contributing to understanding the potential of KANs in computer vision, highlighting both their strengths and areas for further research. Our evaluation [^1] This manuscript is a work in progress, which includes results from an ongoing exploration in this direction. The results and discussion are expected to evolve over time. This independent effort is intended to contributes to our understanding of KANs for visual modeling, and not to provide a yes/no answer to the question ‘if KANs should replace conventional neural modeling techniques for computer vision’. point toward the fact that while KAN-based architectures perform in line with the original claims, it may often be important to employ more complex functions on the network edges to retain the performance advantage of KANs on more complex visual data.
@@ I Introduction
Several implementations and research efforts are currently being made to explore the potential of KANs in various domains. The Convolutional-KANs project [24] extends the idea of KANs to convolutional layers, replacing the classic linear transformation of the convolution with learnable non-linear activations in each pixel. The Vision-KAN repository [25] explores the possibility of MLPs with KANs in Vision Transformers. The KAN-GPT project [26] implements Generative Pre-trained Transformers (GPTs) using KANs for language modeling tasks, demonstrating the potential of KANs in natural language processing applications. The KAN-GPT-2 repository [27] trains small GPT-2 style models using KANs instead of MLPs. The KANeRF project [28] integrates KANs into Neural Radiance Fields (NeRF) for view synthesis tasks. The KAN-UNet repository [29] implements a U-Net architecture with Kolmogorov-Arnold Convolutions for image segmentation tasks. Similarly, the kansformers project [30] explores the use of KANs in Transformer architectures, replacing the traditional linear layers with KAN layers.
Although there is currently a large interest in developing implementation repositories providing various KAN variants pertinent to visual modeling, a formal scientific effort focusing on the suitability of KANs in general and the emerging repositories in particular for visual modeling tasks is still missing. This work is aimed at partially addressing this gap as a preliminary effort focusing on providing independent empirical evidence of KANs’ suitability for fundamental vision task. We replicate KAN-based visual modeling ideas to evaluate their performance in terms of accuracy, training efficiency, and model parameters while experimenting with standard datasets. Our experiments are expected to provide insights into the strengths and limitations of KANs in handling complex visual data. The study aims at helping understand answers to the following questions:
@@ II-E Wavelet KANs (WavKANs)
WavKANConvs provide a powerful extension to KANs, incorporating wavelet-based frequency analysis into the adaptive filtering process of convolutional networks. This combination enables the extraction of complex patterns with enhanced localization in both time and frequency domains, potentially leading to superior performance in various visual tasks.
II-F KANs in Segmentation
Segmentation refers to partitioning an image into meaningful segments at the pixel level, to distinguish different objects or regions. While KANs have shown promise in image recognition tasks, their application to more complex tasks like segmentation has hardly been explored. The U-Net architecture [35] is known for its efficacy in image segmentation and other pixel-wise prediction tasks. It provides an excellent test bed for analyzing KANs, specifically UKAN [36] for segmentation performance. The motivation behind adapting KANs to segmentation tasks stems from the potential of Kolmogorov-Arnold Networks to learn complex, nonlinear mappings more efficiently than traditional neural networks. By incorporating KAN into the U-Net architecture, the learnable functions on the edges are explored.
II-F1 UKAN-Architecture
To investigate the applicability of KANs in segmentation, an adaptation of the U-Net architecture with KAN layers integrated into the convolutional and deconvolutional blocks is adapted from [36], [37].
Figure 4: (A) Overall UKAN Architecture: The UKAN architecture comprises an encoder (Down Blocks), a bottleneck, and a decoder (Up Blocks) pathway. Each block in the encoder and decoder contains UKAN convolutional layers, which are connected by skip connections (dashed lines) to corresponding layers across the encoder and decoder. The input image is progressively downsampled and then upsampled to produce the segmentation output. (B) UKANConv Block: This block, used within the Down and Up blocks, performs two consecutive KConv KAN operations, each followed by Batch Normalization and ReLU activation, which together transform the input feature maps. (C) Conceptual KConvKAN: The KConv KAN operation involves applying learnable spline functions (denoted by $\phi$) to the input features $x_{1}$ and $x_{2}$, producing output features $o_{1}$ and $o_{2}$.
The structure of the network is provided as:
  • Encoder: The encoder in UKAN-UNet follows the typical U-Net design, consisting of convolutional blocks followed by max-pooling layers. However, instead of standard convolutions, we use UKAN layers. These layers replace the conventional linear transformations with learnable, nonlinear spline functions on the edges
  • Bottleneck: The bottleneck, which connects the encoder and decoder, also employs UKAN layers, ensuring that the most abstracted feature maps benefit from the enhanced representational power of KANs.
  • Decoder: In the decoder, UKAN layers are used in the upsampling blocks, allowing the network to reconstruct the spatial details with greater accuracy. The skip connections between corresponding encoder and decoder layers ensure that the high-resolution features are retained, further improved by the flexibility of UKAN.
II-F2 Mathematical Formulation
Let $X$ represent the input image, and $Y$ the output segmentation map. In a traditional U-Net, the convolutional operations can be represented as:
$Y=f_{\text{conv}}(X)$
where $f_{\text{conv}}$ denotes the convolutional layers.
In UKAN-UNet, the convolutional operations are replaced by UKAN layers:
$Y=f_{\text{ukan}}(X)$
where $f_{\text{ukan}}$ represents the UKAN layers, defined as:
$f_{\text{ukan}}(X)=\sum_{m,n}\phi_{m,n}(X_{i+m,j+n})$
Here, $\phi_{m,n}$ are the learnable functions parameterized as splines, allowing the network to adapt more flexibly to the input data.
@@ III Experiments
This section details the experimental setup and procedures used to evaluate the performance of networks on MNIST and CIFAR-10 datasets for classification and Camvid dataset for segmentation. The goal of these experiments is to assess the suitability of KANs for the fundamental computer vision task of recognition, focusing on aspects such as accuracy, model size, training time, and parameter efficiency. The architectural overview and the internal visualizations of a KConvKAN used in our experiments are illustrated in Fig. 3 for a clear understanding. The shown KConvKAN in Fig. 3(a) depicts the flow of data through the network that consists of two KAN Convolutional (KANConv) layers, each followed by a MaxPool2D layer, then a Flatten layer, and finally a KAN Linear layer that maps the flattened output of the convolutional layers to the final classification output via a log-softmax function. Fig.  3(b) provides insights into the feature maps and learned spline weights at different stages of the KConvKAN when processing a sample image. The feature map progression is shown after each KANConv and MaxPool2D layer. The 3D spline weights and heat map representation of spline weights in the KAN Linear layer are also visualized. The visualization is meant to demonstrate the transformation of the input through the network layers for understanding.
III-A KAN in Classification
III-A1 Experimental Setup
@@ III-A4 Results
III-B KAN in Segmentation
III-B1 Experimental Setup
The computation experimental setup is same as outlined in the classification section. The UKAN-UNet was implemented using the PyTorch framework. We initialized the UKAN layers with a grid size of 5 and a spline order of 3. The training was performed using an Adam optimizer with an initial learning rate of 0.001, and the learning rate was decayed exponentially with a gamma of 0.8. The model was trained for 1000 epochs with a batch size of 16 on an NVIDIA A100 GPU.
III-B2 Dataset
CamVid: The CamVid dataset, or Cambridge-driving Labeled Video Database, is a widely used benchmark in the field of semantic segmentation, particularly for autonomous driving applications [39]. This dataset comprises over 700 images extracted from five high-resolution video sequences captured at 30 frames per second using a camera mounted on a vehicle’s dashboard. These sequences were then downsampled to provide labeled still frames at 1 frame per second, resulting in a total of 701 images. Each image is annotated with pixel-level labels corresponding to 32 distinct semantic classes, including categories such as sky, building, road, tree, and pedestrian, among others [40]. For practical purposes, these classes are often condensed into 11 primary categories to facilitate research and model training. The CamVid dataset is particularly valuable for its high-quality annotations and the diversity of urban driving scenes it represents, making it an essential resource for developing and evaluating semantic segmentation algorithms in dynamic environments [41].
III-B3 Results
The performance of UKAN-UNet was evaluated using standard segmentation metrics, including pixel accuracy, mean Intersection over Union (mIoU), and Dice coefficient.
TABLE VII: Test Loss Values for UKAN and UNet at 500 and 1000 Epochs On CamVid Dataset
Metric 500 Epochs - UKAN 500 Epochs - UNet 1000 Epochs - UKAN 1000 Epochs - UNet
Test Loss (Memory) 0.9550 1.0487 0.8443 0.8991
Test Loss (Saved) 0.9800 1.0401 0.8316 0.8872
Test Loss (Final Evaluation) 0.9758 1.0474 0.8296 0.8637
TABLE VIII: Segmentation Results on the CamVid Dataset
Metric UKAN-UNet U-Net
Pixel Accuracy 67.3% 64.5%
mIoU 57.8% 54.2%
Dice Coefficient 64.7% 61.4%
  • Accuracy: UKAN-UNet achieved a pixel accuracy of 67.3% on the CamVid dataset, compared to 64.5% for the baseline U-Net.
  • mIoU: The mean IoU improved from 54.2% with U-Net to 57.8% with UKAN-UNet, indicating better segmentation across all classes.
  • Dice Coefficient: UKAN-UNet also showed an improvement in the Dice coefficient, overall.
III-C Claims Validation
@@
III-C1 Claim-1: Accuracy
The original work on KANs [20] identifies that smaller KAN models can match or exceed the accuracy of larger MLPs in function fitting tasks. We can analyze this claim through our experiments on MNIST, CIFAR10 and CamVid datasets. Our preliminary results show that KANs show competitive results with ConvNets both in terms of classification and segmentation.
III-C2 Claim-2: Computational Efficiency
@@
III-C3 Claim-3: Scalability
According to the KAN paper, these networks exhibit faster neural scaling laws than MLPs. However, when scaling up the networks to accommodate the high-dimensional data from CIFAR10, KANs required significant tuning of spline parameters to maintain performance, which adds to the training complexity. The adaptation of KANs in UNet architecture provides preliminary support to the scalability claim, suggesting that practical scalability is achievable and contingent on the ability to manage increasing model complexity efficiently.
@@ IV Conclusion
This work critically evaluates the efficacy of Kolmogorov-Arnold networks in the basic computer vision task of image recognition, contrasting their performance with conventional networks across MNIST and CIFAR-10 datasets for classification and CamVid dataset for segmentation. This work is a continuous effort that will be updated as more data and results are available. Our current findings reveal that while KANs do offer variation in the traditional neural network architectures, their superior accuracy and scalability are clearly evident for classification and segmentation datasets. The distinctive architecture of KANs, which integrates learnable spline functions directly onto the edges, provides an innovative approach to neural network design and opens new avenues for enhancing model interpretability and efficiency. The practical application of KANs in complex visual tasks require further optimization and exploration to fully harness their theoretical advantages. The research community is already engaged in leveraging KANs for vision or related tasks. Future work should focus on improving KAN-based architecture for scalability and generalization in diverse application scenarios within computer vision. Overall, the comparative analysis indicates that WavKAN’s utilization of wavelet functions not only preserves the strengths of conventional KANs on simpler datasets like MNIST but also extends their applicability and performance on more demanding datasets. It does support the argument that investing in KAN-based vision models is worthwhile. With the large interest of the vision and non-vision community in KANs, we can expect to see interesting developments in KANs in terms of computer vision in the near future.